OpenStudy (anonymous):

2 cos 2θ − 1 = 0 find all solutions of the equation and Find the solutions in the interval [0, 2π)

OpenStudy (anonymous):

x = cos2θ 2x-1=0 x must equal 1/2 cos2θ= 1/2 cos(2*pi/6)=1/2 cos(2*-pi/6) = 1/2

OpenStudy (anonymous):

theta = pi/6 and -pi/6 .. i think.. did it quick